# TFMQ-DM: Temporal Feature Maintenance Quantization for Diffusion Models

### 28 Feb 2024 v2

@@ Abstract
The Diffusion model, a prevalent framework for image generation, encounters significant challenges in terms of broad applicability due to its extended inference times and substantial memory requirements. Efficient Post-training Quantization (PTQ) is pivotal for addressing these issues in traditional models. Different from traditional models, diffusion models heavily depend on the time-step $t$ to achieve satisfactory multi-round denoising. Usually, $t$ from the finite set {1, …, T} is encoded to a temporal feature by a few modules totally irrespective of the sampling data. However, existing PTQ methods do not optimize these modules separately. They adopt inappropriate reconstruction targets and complex calibration methods, resulting in a severe disturbance of the temporal feature and denoising trajectory, as well as a low compression efficiency. To solve these, we propose a Temporal Feature Maintenance Quantization (TFMQ) framework building upon a Temporal Information Block which is just related to the time-step $t$ and unrelated to the sampling data. Powered by the pioneering block design, we devise temporal information aware reconstruction (TIAR) and finite set calibration (FSC) to align the full-precision temporal features in a limited time. Equipped with the framework, we can maintain the most temporal information and ensure the end-to-end generation quality. Extensive experiments on various datasets and diffusion models prove our state-of-the-art results. Remarkably, our quantization approach, for the first time, achieves model performance nearly on par with the full-precision model under 4-bit weight quantization. Additionally, our method incurs almost no extra computational cost and accelerates quantization time by 2.0×2.0$\times$2.0 × on LSUN-Bedrooms $256\times 256$ compared to previous works.
@@ 1 Introduction
Quantization is currently the most widely used method [32, 7, 33, 5, 1] for compressing models by mapping high-precision floating-point numbers to low-precision numbers. Among different quantization methods, Post-training quantization (PTQ) [32, 48, 15] incurs lower overhead and is more user-friendly without the need for retraining or fine-tuning. While PTQ on conventional models has undergone extensive study [32, 24, 48, 7], its application to diffusion models has shown huge performance degradation, especially under low-bit settings. For instance, Q-Diffusion [23] exhibits severe accuracy drop on some datasets [49] under 4-bit quantization. We believe the reason they fail to achieve better results is that they all overlook the sampling data independence and uniqueness of temporal features, which are generated from time-step $t$ through a few modules, used to control the denoising trajectory in diffusion models. Therefore, we observe that temporal feature disturbance significantly impacts model performance in the aforementioned methods.
@@
  • A framework of temporal feature maintenance quantization (TFMQ) is proposed, consisting of temporal information aware reconstruction (TIAR) for weight quantization and finite set calibration (FSC) for activation quantization. Both are based on a Temporal Information Block specially devised for diffusion models.
@@ 4.1 Temporal Feature Disturbance
where $\text{cos}(\cdot)$ denotes cosine similarity, and $\widehat{\mathbf{emb}_{t,i}}$ signifies the temporal feature corresponding to $\mathbf{emb}_{t,i}$ in the quantized model. As illustrated in Fig. 3, quantization induces notable temporal feature errors. We term this phenomenon, characterized by substantial temporal feature errors within diffusion models, as temporal feature disturbance.
@@
Furthermore, as depicted in Fig. 3, we even observe a pronounced temporal information mismatch. Specifically, the temporal feature generated by the quantized model at time-step $t$ exhibits a divergence from that of the full-precision model at the corresponding time-step, Instead, it tends to align more closely with the temporal feature corresponding to $t+\delta_{t}$, importing wrong temporal information from $t+\delta_{t}$.
Figure 2: Temporal feature disturbance. The inflection points serve as indicators of temporal feature errors at different time-steps, and they highlight the significant phenomenon of temporal feature disturbance.
Figure 2: Temporal feature disturbance. The inflection points serve as indicators of temporal feature errors at different time-steps, and they highlight the significant phenomenon of temporal feature disturbance.
Figure 3: Temporal information mismatch. The coordinates of the inflection points on the blue curve can denoted as $(t,t+\delta_{t,i})$. It indicates $\mathbf{emb}_{t+\delta_{t,i},i}$ exhibits the highest similarity with $\widehat{\mathbf{emb}_{t,i}}$.
@@ 4.2 Inducement Analyses
In this section, we explore the two inducements of temporal feature disturbance. For the purpose of clarity, in the subsequent sections, “reconstruction” specifically points to slight weight adjustment for minimal quantization error, while “calibration” specifically refers to activation calibration.
@@ 4.3 Temporal Feature Maintenance
$\mathcal{L}_{TIB}=\sum_{i=0}^{n}\|g_{i}(h(t))-\widehat{g}_{i}(\widehat{h}(t))\|_{F}^{2},$ (10)
@@ 5.1 Implementation details
Evaluation metrics. For each experiment, we evaluate the performance of diffusion models with Fréchet Inception Distance (FID) [12]. In the case of LDM and Stable Diffusion experiments, we also include sFID [41], which better captures spatial relationships than FID. For ImageNet and CIFAR-10 experiments, we additionally provide Inception Score (IS) [41] as a reference metric. Further, in the context of Stable Diffusion experiments, we extend our evaluation to include the compatibility of image-caption pairs, employing the CLIP score [11]. The ViT-B/32 is used as the backbone when computing the CLIP score. To ensure consistency in the reported outcomes, all results are derived from our own implementation or from other papers, where experiments are conducted under conditions consistent with ours. More specifically, in the evaluation process of each experiment, we sample 50k images from DDPM or LDM, or 30k images from Stable-Diffusion. All experiments are conducted utilizing one H800 GPU and implemented with the PyTorch framework [35].
@@ 5.2 Main Results
In the experiments conducted on the LDM, we maintain the same experimental settings as presented in  [39], including the number of steps, variance schedule, and classifier-free guidance scale (denoted by eta and cfg in the following, respectively). As shown in Tab. 2, the FID performance differences relative to the full precision (FP) model are all within 0.7 for all settings. Specifically, on the CelebA-HQ $256\times 256$ dataset, our method exhibits a FID reduction of 6.71 and a sFID reduction of 6.60 in the w4a8 setting compared to the current state-of-the-art (SOTA). It is noticeable that existing methods, whether in 4-bit or 8-bit, show significant performance degradation when compared to the FP model on face datasets like CelebA-HQ $256\times 256$ and FFHQ $256\times 256$, whereas our TFMQ-DM shows almost no performance degradation compared to the FP model. Importantly, our method achieves significant performance improvement on the LSUN-Bedrooms $256\times 256$ in the w4a8 setting, with FID and sFID reductions of 2.26 and 7.51 compared to PTQD [10], respectively. Regarding LDM-8 on LSUN-Churches $256\times 256$, we attribute the moderate improvement, compared to other datasets. We believe that the use of the LDM-8 model with a downsampling factor of 8 may be more quantization-friendly. Existing methods have already achieved satisfactory results on this dataset. Nonetheless, our method still approaches the performance of the FP model more closely compared to existing methods.
Table 3: Quantization results for unconditional image generation with DDIM on CIFAR-10 $32\times 32$.
Methods Bits (W/A) CIFAR-10 $32\times 32$
[…]
Full Prec. 32/32 9.04 4.23
PTQ4DM* [42] 4/32 9.02 5.65
Q-Diffusion† [23] 4/32 8.78 5.08
TDQ [44] 4/32 - -
TFMQ-DM (Ours) 4/32 9.14 (+0.12) 4.73 (-0.35)
PTQ4DM [42] 8/8 9.02 19.59
Q-Diffusion† [23] 8/8 8.89 4.78
TDQ [44] 8/8 8.85 5.99
TFMQ-DM (Ours) 8/8 9.07 (+0.05) 4.24 (-0.54)
PTQ4DM* [42] 4/8 8.93 5.14
Q-Diffusion† [23] 4/8 9.12 4.98
TDQ [44] 4/8 - -
TFMQ-DM (Ours) 4/8 9.13 (+0.01) 4.78 (-0.20)
We also conduct experiments with DDPM on CIFAR-10 $32\times 32$. Due to the lower resolution and simplicity of the images in this dataset, existing methods show minimal performance degradation. As seen in Tab. 3, while the results we obtain may not be as pronounced, we still achieve comprehensive improvements in terms of IS and FID compared to the existing SOTA.
@@ Appendix C Inappropriate Calibration Target
Appendix D Additional Effect of TIAR
@@
Appendix E Inference Cost of TSC
In this section, we assess the inference overhead of our TFMQ-DM method on real hardware, specifically the ${}^{\circledR}$ ${}^{\circledR}$ Gold 6248R Processor. All floating-point and quantized operations are implemented using Intel’s OpenVINO toolkit  [^3] OpenVINO toolkit. As illustrated in Table II, in comparison to the UNet quantized with the built-in w8a8 quantization method in the OpenVINO toolkit, our approach results in a memory overhead of less than $0.076\%$, yielding a 2.38×2.38$\times$2.38 × acceleration compared to the original floating-point model. Moreover, our method introduces less than $0.5\%$ additional latency compared to the built-in w8a8 quantization in the OpenVINO toolkit.
Table II: Inference analysis of Stable Diffusion with 50 denoising time-steps on Intel CPU.
@@ Appendix E Inference Cost of TSC
Appendix F Study of Sampling with Advanced Samplers
Apart from employing the DDIM sampler [45], we also utilize a variant of DDPM [14] called PLMS [27] on the CelebA-HQ $256\times 256$ dataset [18]. This better demonstrates the superiority of our TFMQ-DM framework compared to previous works. From Tab. III, the introduced TFMQ-DM substantially reduces FID and sFID, surpassing PTQD by margins of 12.40 and 7.09, respectively.
Table III: Quantization results for unconditional image generation with PLMS on CelebA-HQ $256\times 256$.
@@ Appendix F Study of Sampling with Advanced Samplers
Additionally, we present experiments performed using the DPM++ solver [30] on LSUN-Churches $256\times 256$ [49]. As illustrated in Tab. IV, our framework consistently outperforms existing methods in terms of performance on this dataset with the DPM++ solver.
Table IV: Quantization results for unconditional image generation with DPM++ on LSUN-Churches $256\times 256$.
@@
Appendix G Comparison of Visualization Results
Within this section, we present random samples derived from full-precision and w4a8 quantized diffusion models with a fixed random seed. These quantized models were created through our TFMQ-DM or previous state-of-the-art methods. The figures below illustrate the obtained results. As shown from Fig. III to Fig. IX, our framework yields results that closely resemble those of the full-precision model, showcasing higher fidelity. Moreover, it excels in finer details, producing superior outcomes in some intricate aspects (zoom in to closely examine the relevant images).
@@ Appendix G Comparison of Visualization Results
Figure IV: Random samples from w4a8 quantized and full-precision LDM-4 on FFHQ $256\times 256$. The resolution of each sample is $256\times 256$.
@@
[…]
Figure V: Random samples from w4a8 quantized and full-precision LDM-8 on LSUN-Churches $256\times 256$. The resolution of each sample is $256\times 256$.
@@
Figure VI: Random samples from w4a8 quantized and full-precision LDM-4 on LSUN-Bedrooms $256\times 256$. The resolution of each sample is $256\times 256$.
@@
Figure VII: Random samples from w4a8 quantized and full-precision LDM-4 on ImageNet $256\times 256$. The resolution of each sample is $256\times 256$.
@@
Figure VIII: Random samples from w4a8 quantized and full-precision DDIM on CIFAR-10 $32\times 32$. The resolution of each sample is $32\times 32$.
[…]
Figure IX: Random samples from w4a8 quantized and full-precision Stable Diffusion. (Left) prompt: A digital illustration of the Babel tower, detailed, trending in artstation, fantasy vivid colors. (Right) prompt: A beautiful castle beside a waterfall in the woods. The resolution of each sample is $512\times 512$.

### 11 Mar 2024 v3

@@ Abstract
The Diffusion model, a prevalent framework for image generation, encounters significant challenges in terms of broad applicability due to its extended inference times and substantial memory requirements. Efficient Post-training Quantization (PTQ) is pivotal for addressing these issues in traditional models. Different from traditional models, diffusion models heavily depend on the time-step $t$ to achieve satisfactory multi-round denoising. Usually, $t$ from the finite set $\{1,\ldots,T\}$ is encoded to a temporal feature by a few modules totally irrespective of the sampling data. However, existing PTQ methods do not optimize these modules separately. They adopt inappropriate reconstruction targets and complex calibration methods, resulting in a severe disturbance of the temporal feature and denoising trajectory, as well as a low compression efficiency. To solve these, we propose a Temporal Feature Maintenance Quantization (TFMQ) framework building upon a Temporal Information Block which is just related to the time-step $t$ and unrelated to the sampling data. Powered by the pioneering block design, we devise temporal information aware reconstruction (TIAR) and finite set calibration (FSC) to align the full-precision temporal features in a limited time. Equipped with the framework, we can maintain the most temporal information and ensure the end-to-end generation quality. Extensive experiments on various datasets and diffusion models prove our state-of-the-art results. Remarkably, our quantization approach, for the first time, achieves model performance nearly on par with the full-precision model under 4-bit weight quantization. Additionally, our method incurs almost no extra computational cost and accelerates quantization time by 2.0×2.0$\times$2.0 × on LSUN-Bedrooms $256\times 256$ compared to previous works. Our code is publicly available at https://github.com/ModelTC/TFMQ-DM.
@@ 1 Introduction
Quantization is currently the most widely used method [32, 7, 33, 5, 1] for compressing models by mapping high-precision floating-point numbers to low-precision numbers. Among different quantization methods, Post-training quantization (PTQ) [32, 48, 15] incurs lower overhead and is more user-friendly without the need for retraining or fine-tuning. While PTQ on conventional models has undergone extensive study [32, 24, 48, 7], its application to diffusion models has shown huge performance degradation, especially under low-bit settings. For instance, Q-Diffusion [23] exhibits severe accuracy drop on some datasets [49] under 4-bit quantization. We believe the reason they fail to achieve better results is that they all overlook the sampling data independence and uniqueness of temporal features, which are generated from time-step $t$ through a few modules, used to control the denoising trajectory in diffusion models. Furthermore, we observe that temporal feature disturbance significantly impacts model performance in the aforementioned methods.
@@
  • An advanced framework (TFMQ-DM) is proposed, consisting of temporal information aware reconstruction (TIAR) for weight quantization and finite set calibration (FSC) for activation quantization. Both are based on a Temporal Information Block specially devised for diffusion models.
@@ 4.1 Temporal Feature Disturbance
where $\text{cos}(\cdot)$ denotes cosine similarity, and $\widehat{\mathbf{emb}_{t,i}}$ signifies the temporal feature corresponding to $\mathbf{emb}_{t,i}$ in the quantized model. As illustrated in Fig. 2 (Left), quantization induces notable temporal feature errors. We term this phenomenon, characterized by substantial temporal feature errors within diffusion models, as temporal feature disturbance.
@@
Furthermore, as depicted in Fig. 2 (Right), we even observe a pronounced temporal information mismatch. Specifically, the temporal feature generated by the quantized model at time-step $t$ exhibits a divergence from that of the full-precision model at the corresponding time-step, Instead, it tends to align more closely with the temporal feature corresponding to $t+\delta_{t}$, importing wrong temporal information from $t+\delta_{t}$.
Figure 2: (Left) Temporal feature disturbance. The inflection points serve as indicators of temporal feature errors at different time-steps, and they highlight the significant phenomenon of temporal feature disturbance. (Right) Temporal information mismatch. The coordinates of the inflection points on the blue curve can denoted as $(t,t+\delta_{t,i})$. It indicates $\mathbf{emb}_{t+\delta_{t,i},i}$ exhibits the highest similarity with $\widehat{\mathbf{emb}_{t,i}}$.
@@ 4.2 Inducement Analyses
In this section, we explore the two inducements of temporal feature disturbance. For the purpose of clarity, in the subsequent sections, “reconstruction" specifically points to slight weight adjustment for minimal quantization error, while “calibration" specifically refers to activation calibration.
@@ 4.3 Temporal Feature Maintenance
$\mathcal{L}_{TIAR}=\sum_{i=0}^{n}\|g_{i}(h(t))-\widehat{g}_{i}(\widehat{h}(t))\|_{F}^{2},$ (10)
@@ 5.1 Implementation details
Evaluation metrics. For each experiment, we evaluate the performance of diffusion models with Fréchet Inception Distance (FID) [12]. In the case of LDM and Stable Diffusion experiments, we also include sFID [41], which better captures spatial relationships than FID. For ImageNet and CIFAR-10 experiments, we additionally provide Inception Score (IS) [41] as a reference metric. Further, in the context of Stable Diffusion experiments, we extend our evaluation to include the compatibility of image-caption pairs, employing the CLIP score [11]. The ViT-B/32 is used as the backbone when computing the CLIP score. To ensure consistency in the reported outcomes, all results are derived from our implementation or from other papers, where experiments are conducted under conditions consistent with ours. More specifically, in the evaluation process of each experiment, we sample 50k images from DDPM or LDM, or 30k images from Stable-Diffusion. All experiments are conducted utilizing one H800 GPU and implemented with the PyTorch framework [35].
@@ 5.2 Main Results
In the experiments conducted on the LDM, we maintain the same experimental settings as presented in  [39], including the number of steps, variance schedule, and classifier-free guidance scale (denoted by eta and cfg in the following, respectively). As shown in Tab. 2, the FID performance differences relative to the full precision (FP) model are all within 0.7 for all settings. Specifically, on the CelebA-HQ $256\times 256$ dataset, our method exhibits a FID reduction of 6.71 and a sFID reduction of 6.60 in the w4a8 setting compared to the current state-of-the-art (SOTA). It is noticeable that existing methods, whether in 4-bit or 8-bit, show significant performance degradation when compared to the FP model on face datasets like CelebA-HQ $256\times 256$ and FFHQ $256\times 256$, whereas our TFMQ-DM shows almost no performance degradation compared to the FP model. Importantly, our method achieves significant performance improvement on the LSUN-Bedrooms $256\times 256$ in the w4a8 setting, with FID and sFID reductions of 2.26 and 7.51 compared to PTQD [10], respectively. Regarding LDM-8 on LSUN-Churches $256\times 256$, we attribute the moderate improvement, compared to other datasets. We believe that the use of the LDM-8 model with a downsampling factor of 8 may be more quantization-friendly. Existing methods have already achieved satisfactory results on this dataset. Nonetheless, our method still approaches the performance of the FP model more closely compared to existing methods. Besides the experiments on LDM, We have also conducted experiments with DDPM on CIFAR-10 $32\times 32$, which can be found in the appendix.
@@ 6 Conclusion & Discussions of Limitations
However, besides temporal features, we have also found the textual features introduced in Stable Diffusion have a physical meaning and influence the generation effect. Nevertheless, these textual features are not taken into consideration in current methods. Secondly, temporal feature maintenance is suitable for both PTQ and QAT scenarios. But in this study, our main focus is on the PTQ setting and achieving 4-bit quantization. The extension to the QAT setting should be conducted in the future to achieve lower-bit quantization and further performance enhancements.
@@ Appendix C Inappropriate Calibration Target
Appendix D Unconditional Image Generation on CIFAR-10
In this section, we conduct more experiments for unconditional image generation on CIFAR-10 $32\times 32$. As shown in Fig. II, our methods still achieve comprehensive improvements in terms of IS and FID compared to the existing SOTA. However, due to the lower resolution and simplicity of the images in this dataset, existing methods show minimal performance degradation, so the results we obtain may not be as pronounced.
Table II: Quantization results for unconditional image generation with DDIM on CIFAR-10 $32\times 32$.
Methods Bits (W/A) CIFAR-10 $32\times 32$
[…]
Full Prec. 32/32 9.04 4.23
PTQ4DM* [42] 4/32 9.02 5.65
Q-Diffusion† [23] 4/32 8.78 5.08
TDQ [44] 4/32 - -
TFMQ-DM (Ours) 4/32 9.14 (+0.12) 4.73 (-0.35)
PTQ4DM [42] 8/8 9.02 19.59
Q-Diffusion† [23] 8/8 8.89 4.78
TDQ [44] 8/8 8.85 5.99
TFMQ-DM (Ours) 8/8 9.07 (+0.05) 4.24 (-0.54)
PTQ4DM* [42] 4/8 8.93 5.14
Q-Diffusion† [23] 4/8 9.12 4.98
TDQ [44] 4/8 - -
TFMQ-DM (Ours) 4/8 9.13 (+0.01) 4.78 (-0.20)
Appendix E Additional Effect of TIAR
@@
Appendix F Inference Cost of TSC
In this section, we assess the inference overhead of our TFMQ-DM method on real hardware, specifically the ${}^{\circledR}$ ${}^{\circledR}$ Gold 6248R Processor. All floating-point and quantized operations are implemented using Intel’s OpenVINO toolkit  [^3] OpenVINO toolkit. As illustrated in Table III, in comparison to the UNet quantized with the built-in w8a8 quantization method in the OpenVINO toolkit, our approach results in a memory overhead of less than $0.076\%$, yielding a 2.38×2.38$\times$2.38 × acceleration compared to the original floating-point model. Moreover, our method introduces less than $0.5\%$ additional latency compared to the built-in w8a8 quantization in the OpenVINO toolkit.
Table III: Inference analysis of Stable Diffusion with 50 denoising time-steps on Intel CPU.
@@ Appendix F Inference Cost of TSC
Appendix G Study of Sampling with Advanced Samplers
Apart from employing the DDIM sampler [45], we also utilize a variant of DDPM [14] called PLMS [27] on the CelebA-HQ $256\times 256$ dataset [18]. This better demonstrates the superiority of our TFMQ-DM framework compared to previous works. From Tab. IV, the introduced TFMQ-DM substantially reduces FID and sFID, surpassing PTQD by margins of 12.40 and 7.09, respectively.
Table IV: Quantization results for unconditional image generation with PLMS on CelebA-HQ $256\times 256$.
@@ Appendix G Study of Sampling with Advanced Samplers
Additionally, we present experiments performed using the DPM++ solver [30] on LSUN-Churches $256\times 256$ [49]. As illustrated in Tab. V, our framework consistently outperforms existing methods in terms of performance on this dataset with the DPM++ solver.
Table V: Quantization results for unconditional image generation with DPM++ on LSUN-Churches $256\times 256$.
@@
Appendix H Comparison of Visualization Results
Within this section, we present random samples derived from full-precision and w4a8 quantized diffusion models with a fixed random seed. These quantized models were created through our TFMQ-DM or previous state-of-the-art methods. The figures below illustrate the obtained results. As shown from Fig. III to Fig. VIII, our framework yields results that closely resemble those of the full-precision model, showcasing higher fidelity. Moreover, it excels in finer details, producing superior outcomes in some intricate aspects (zoom in to closely examine the relevant images).
@@ Appendix H Comparison of Visualization Results
Figure IV: Random samples from w4a8 quantized and full-precision LDM-8 on LSUN-Churches $256\times 256$. The resolution of each sample is $256\times 256$.
@@
Figure V: Random samples from w4a8 quantized and full-precision LDM-4 on LSUN-Bedrooms $256\times 256$. The resolution of each sample is $256\times 256$.
@@
Figure VI: Random samples from w4a8 quantized and full-precision LDM-4 on ImageNet $256\times 256$. The resolution of each sample is $256\times 256$.
@@
Figure VII: Random samples from w4a8 quantized and full-precision DDIM on CIFAR-10 $32\times 32$. The resolution of each sample is $32\times 32$.
@@
Figure VIII: Random samples from w4a8 quantized and full-precision Stable Diffusion. (Left) prompt: A digital illustration of the Babel tower, detailed, trending in artstation, fantasy vivid colors. (Right) prompt: A beautiful castle beside a waterfall in the woods. The resolution of each sample is $512\times 512$.